OpenStudy (anonymous):

Please help (: Thank you ! For questions 1 & 2, add or subtract? 1. (8x - 2y) + (3x - 4y) a) 11x + 8y b) 4x - y c) 11x - 6y d) 5x + 2y 2. (y + 3d) - (-9y + 7d) a) -10y + 4d b) -8y + 10d c) 8y - 6d d) 10y - 4d

OpenStudy (anonymous):

1 is a simple gathering of terms. Be careful on 2. You are subtracting EVERYTHING in the parentheses.

OpenStudy (johnweldon1993):

1) combine like terms 2) distribute that '-' sign into the parenthesis first...then add the 2 equations

OpenStudy (anonymous):

Yeah, I still remember from when you helped last time (: But on the first one, since its a '-' sign do I subtract ?

OpenStudy (anonymous):

Or is it visa versa, meaning since its minus, I add?

OpenStudy (reemii):

you have to collect the coefficients , for example, related to the variable "y". you must add them up. the coefficients are (-2) and (-4). Sum: (-2)+(-4) = -2-4 = -6.

OpenStudy (reemii):

this means that in the answer, the coefficient of the variable y is -6. it's done like that: \(\text{stuff} -2y + \text{stuff} - 4y = \text{stuff} + \text{stuff} +(-2-4)y = \text{stuff}+\text{stuff}-6y\).

OpenStudy (anonymous):

So the answer is 11x - 6y? Or did I do that wrong?

OpenStudy (reemii):

you're right.

OpenStudy (anonymous):

Thank you :)